OpenStudy (anonymous):

The fraction (x/y) + x divided by (1/y) + 1 is equal to? The answer is x but how do you get to that result?

OpenStudy (anonymous):

The x is being added to x/y and the 1 is being added to 1/y

OpenStudy (anonymous):

Note that:\[\frac{x}{y}+x=\left(\frac{1}{y}+1\right)x\]

OpenStudy (pfenn1):

\[\left( \frac{x}{y} + x \right) \div \left( \frac{1}{y} + 1 \right)=\]

OpenStudy (anonymous):

\[\frac{\left(\frac{x}{y}+x\right)}{\left(\frac{1}{y}+y\right)}=\frac{\left(\frac{1}{y}+1\right)x}{\left(\frac{1}{y}+y\right)}\]is it clear what happens after this?

OpenStudy (anonymous):

Wouldn't the setup be that of a complex fraction?

OpenStudy (anonymous):

It needs to be simplified and when simplified completely the answer is x, I don't know how to get to x though

OpenStudy (anonymous):

But how did you get x outside of the parenthesis? I understand how you crossed that out, but how are you getting x out of the parenthesis and getting it alone?